# Addressing Data Heterogeneity in Federated Learning with Adaptive Normalization-Free Feature Recalibration

### v1

@@ 4.1 Experimental Settings
Datasets. We evaluate our approach on five classification datasets, including Fed-ISIC2019 (Ogier du Terrail et al., 2022) containing dermoscopy images from 6 centers across 8 classes where label distribution skew and heavy quantity skew is present; FedChest, a novel chest X-Ray multi-label dataset with 4 clients and 8 labels with label distribution skew and covariate shift; a partitioning of CIFAR-10 (Krizhevsky et al., 2009) which simulates heavy label distribution skew across 8 clients using the Kolmogorov-Smirnov (KS) ‘split-2’ as presented in Qu et al. (2022); CelebA (Liu et al., 2015) from the LEAF suite (Caldas et al., 2018) which is a binary classification task in a cross-device setting with a large number of clients, covariate shift and high quantity skew; and FedPathology, a colorectal cancer pathology slide dataset with 9 classes derived from Kather et al. (2019), featuring challenging concept drift as the images, which we do not color-normalize, were produced using two different staining protocols. FedChest contains images from PadChest (Bustos et al., 2020), CXR-14 (Wang et al., 2017) and CheXpert Irvin et al. (2019), which present one or more of 8 common disease labels. For FedPathology which we use for DP training in Section 4.3, we use Dirichlet distribution sampling (Hsu et al., 2019) with to simulate a moderate label distribution skew and partition the data among 3 clients. Each task covers a different aspect of the multi-faceted problem of data heterogeneity in FL, including different domains and sources of heterogeneity, to provide a robust test bed. Additional dataset details are presented in Appendix A.1, including instructions to replicate FedChest in B.1.
@@
Compared models. We compare our approach with a typical ResNet (utilizing BN), a ResNet where BN is replaced by GN, a SE-ResNet (Hu et al., 2018), and a NF-ResNet. This selection isolates the effects of our architectural changes compared to using BN, using its popular substitution GN, and using weight standardization and CA separately. We choose a depth of 50 layers for all models to balance performance with computational expense. All models are pre-trained on ImageNet (Russakovsky et al., 2015) using the timm (Wightman, 2019) open-source library. ANFR follows the structure of NF-ResNet, with the addition of CA blocks in the same position as SE-ResNet. Except for Section 4.4, we employ Squeeze-and-Excitation (Hu et al., 2018) as the attention mechanism. Additional model details are provided in Appendix A.3.
@@
Evaluated methods. We use 4 global FL (GFL) and 2 personalized FL (pFL) aggregation methods as axes of comparison for the models, each representing a different approach to model aggregation: the seminal FedAvg (McMahan et al., 2017) algorithm, FedProx (Li et al., 2020a), which adds a proximal loss term to mitigate drift between local and global weights, SCAFFOLD (Karimireddy et al., 2020), which corrects client drift by using control variates to steer local updates towards the global model, FedAdam (Reddi et al., 2021), which decouples server-side and client-side optimization and employs the Adam optimizer (Kingma & Ba, 2017) at the server for model aggregation, FedBN (Li et al., 2021) which accommodates data heterogeneity by allowing clients to maintain their personal batch statistics, and by construction is only applicable to models with BN layers, and FedPer (Arivazhagan et al., 2019) which personalizes the FL process by keeping the weights of the classifier head private to each client. We note our proposal is an architectural one which is aggregation method-agnostic, thus we selected these widely known aggregation methods to represent a spectrum of strategies, from standard averaging to methods addressing client drift and personalization. This provides a robust comparison concentrated on the model architectures.
@@
Evaluation metrics. For Fed-ISIC2019, we report the average balanced accuracy due to heavy class-imbalance as in (Ogier du Terrail et al., 2022). For FedChest, a multi-label classification task with imbalanced classes, we report the mean AUROC on the held-out test set to account for varying class distributions, in addition to more extensive results in Appendix B.2. We report the average accuracy for the other 3 datasets. In pFL settings, the objective is providing good in-federation models. Consequently, we report the average metrics of the best local models, as suggested in (Zhang et al., 2023).
@@
Implementation Details. We choose training hyper-parameters separately for each dataset by tuning the vanilla ResNet using the ranges detailed in Appendix A.2 and then re-using the same parameters for all models. In Fed-ISIC2019 clients use Adam with a learning rate of 5e-4 and a batch size of 64 to train for 80 rounds of 200 steps. In FedChest clients use Adam with a learning rate of 5e-4 and a batch size of 128 to train for 20 rounds of 200 steps. For DP-training in FedPathology, we set the probability of information leakage to , as is common, the noise multiplier to 1.1, the gradient max norm to 1.0, and train for 25 rounds, which is the point where the models have expended a privacy budget of . For CelebA and CIFAR-10 we follow the settings of Qu et al. (2022); Pieri et al. (2023) which were tuned by the authors. All experiments are run in a simulated FL environment with NVFLARE (Roth et al., 2022) and PyTorch (Paszke et al., 2019), using 2 NVIDIA A100 GPUs for training. We report the mean and standard deviation across 3 seeds.
4.2 Performance Analysis and Comparison
@@
GFL scenario. Average results for all datasets, models, and GFL aggregation methods are presented in table 1. First, we observe that GN does not consistently outperform the vanilla ResNet, supporting our pursuit of a more reliable alternative. For instance, GN is outperformed by BN in half of the tested aggregation methods on Fed-ISIC2019 and FedChest. Second, the sub-optimality of CA operating on BN-normalized features is evident, as the SE model frequently performs worse than BN-ResNet, notably across all aggregation methods on FedChest. NF-ResNet shows strong performance across all tasks and methods, confirming the potential of replacing activation normalization with weight standardization in FL. However, our proposed ANFR model consistently outperforms NF-ResNet, often by a considerable margin. For example, on Fed-ISIC2019 with SCAFFOLD, ANFR surpasses NF-ResNet’s mean balanced accuracy by more than 3%. For the FedChest dataset, we employ a large batch size of 128 to maximize the probability that all classes are represented in each batch, following best practices. This choice incidentally biases the setting in favor of BN models since larger batches reduce inconsistent averaging of mini-batch statistics and BN parameters. Despite this bias, ANFR emerges as the top-performing model across aggregation methods. These results indicate that integrating CA with SWS networks provides significant performance gains, suggesting that channel attention is a crucial component in designing effective FL models.
@@
pFL scenario. Table 2 presents the results for pFL scenarios on Fed-ISIC2019 and FedChest. In FedChest, where images are grayscale and we use a large batch size, FedBN and FedPer are virtually equal: BN-ResNet achieves an AUROC of 83.38% with FedBN and 83.39% with FedPer, indicating that the estimated BN statistics of the local models closely match the true ones. GN-ResNet attains 83.73% with FedPer, slightly outperforming BN-ResNet, but ANFR with FedPer is the most performant option across both aggregation methods, yielding a mean AUROC of 83.8%. Conversely, under the severe label and quantity skew on Fed-ISIC2019, employing FedBN improves performance over FedPer for models employing BN. ANFR achieves the highest balanced accuracy of 84.94% nonetheless. Notably, GN performs worse than BN on Fed-ISIC2019, and the ineffectiveness of combining BN and CA is further evidenced, as SE-ResNet is outperformed by BN-ResNet in all scenarios. These findings demonstrate that adopting ANFR enhances performance across both datasets, leading to the best overall models. Unlike the trade-offs observed with BN-FedBN and GN-FedPer combinations, ANFR consistently outperforms other architectures across varying levels of data heterogeneity.
@@
Cross-device experiments on CelebA. Table 3 presents the results of our models on the cross-device setting of CelebA, which contains 200,288 samples across 9,343 clients. While the binary classification task is relatively straightforward for individual clients, it poses challenges at the server level due to the vast number of clients and significant quantity and class skews—some clients have only a few samples or labels from a single class. We observe that ANFR outperforms the baseline models, demonstrating its adaptability across diverse FL scenarios.
@@ 4.3 Sample-level Differentially Private Training
In privacy-preserving scenarios involving differential privacy (DP) BN cannot be used because calculating mini-batch statistics violates privacy-preservation rules, and is customarily replaced by GN. We demonstrate the utility of ANFR compared to GN in such settings using the FedPathology set-up described in Section 4.1. We first train with strict sample-level privacy guarantees, employing a privacy budget of = 1, followed by training without privacy constraints ( = ) to illustrate the privacy/utility trade-off of each model.
@@
From the results presented in Table 4, we observe that with an unrestricted privacy budget, GN and ANFR perform comparably. However, when a strict budget is enforced GN suffers a sharp performance decrease of 17%, as expected following previous research (Klause et al., 2022), whereas ANFR’s average accuracy is reduced by only 3%. ANFR’s robustness under DP may be attributed to its reliance on weight standardization, which has been shown to benefit from additional regularization (Brock et al., 2021b; Zhuang & Lyu, 2024) such as that provided by DP-SGD’s gradient clipping and gradient noising. Our experiments show DP training induces a regularization effect that disproportionately benefits NF models like ANFR, an observation also reported by De et al. (2022). These findings make ANFR a promising candidate for furthering development and adoption of DP training in FL, thereby enhancing the privacy of source data contributors, such as patients.
@@ 4.4 Attention Mechanism Comparison
Next, we investigate the impact of different attention mechanisms on performance. We compare the SE module used in previous sections with ECA (Wang et al., 2020), and CBAM (Woo et al., 2018). ECA replaces SE’s fully-connected layers with a more efficient 1-D convolution to capture local cross-channel interactions. CBAM combines channel and spatial attention and utilizes both max and average pooling to extract channel representations. From Table 5 we observe that even the lowest-performing module on each dataset outperforms all baseline models from Tables 1 and 3, proving the robustness of our approach. No single mechanism consistently performs best, making further exploration of attention module designs an interesting avenue for future work.
@@
4.5 Qualitative Localization Performance Comparison
Finally, we assess the localization capability of each architecture after FL training with the best aggregation method on FedChest, SCAFFOLD. We compare the bounding box annotations provided by Wang et al. (2017) with Grad-CAM++ (Chattopadhay et al., 2018) heatmaps generated for samples labeled Atelectasis or Pneumonia from the FedChest test set. Figure 4 shows that ANFR’s heatmaps closely align with the annotated bounding boxes, indicating better localization of pathologies compared to baselines. This improved localization aids model interpretability, which is crucial in areas like medical imaging.

### v2

@@ 4.1 Experimental Settings
Datasets. We evaluate our approach on five classification datasets, including Fed-ISIC2019 (Ogier du Terrail et al., 2022) containing dermoscopy images from 6 centers with 8 classes where label distribution skew and heavy quantity skew is present; FedChest, a novel chest X-Ray multi-label dataset with 4 clients and 8 labels with label distribution skew and covariate shift; a highly non-IID partitioning of CIFAR-10 (Krizhevsky et al., 2009) which simulates heavy label distribution skew across 5 clients using the Kolmogorov-Smirnov (KS) ‘split-2’ as presented in Qu et al. (2022); CelebA (Liu et al., 2015) from the LEAF suite (Caldas et al., 2018), a binary classification task in a cross-device setting with a large number of clients, covariate shift and high quantity skew; and FedPathology, a colorectal cancer pathology slide dataset with 9 classes derived from Kather et al. (2019), featuring challenging concept drift as the images, which we do not color-normalize, were produced using two different staining protocols. FedChest contains images from PadChest (Bustos et al., 2020), CXR-14 (Wang et al., 2017) and CheXpert Irvin et al. (2019), which present one or more of 8 common disease labels. For FedPathology, used for DP training in Section 4.4, Dirichlet distribution sampling (Hsu et al., 2019) with is used to simulate a moderate label distribution skew and partition the data to 3 clients. Each task covers a different aspect of the multi-faceted problem of data heterogeneity in FL, including different domains and sources of heterogeneity, to provide a robust test bed. More details are presented in Appendix A.1, including instructions to replicate FedChest in C.1.
@@
Compared models. In Sections 4.2-4.7, we compare ANFR with a typical ResNet (utilizing BN), a ResNet where BN is replaced by GN, a SE-ResNet (Hu et al., 2018), and a NF-ResNet. This selection isolates the effects of our architectural changes compared to using BN, using its popular substitution GN, and using weight standardization and CA separately. We choose a depth of 50 layers for all models to balance performance with computational expense. In Section 4.8, we broaden our investigation to include models from different families. All models used in Section 4 are pre-trained on ImageNet (Russakovsky et al., 2015) using timm (Wightman, 2019), but additional experiments with randomly initialized models are presented in Appendix B.2. ANFR follows the structure of NF-ResNet, with the addition of CA blocks in the same position as SE-ResNet. Except for Section 4.5, we employ Squeeze-and-Excitation (Hu et al., 2018) as the attention mechanism. Additional model and computational overhead details are provided in Appendix A.3.
@@
Evaluated methods. We use 4 global FL (GFL) and 3 personalized FL (pFL) aggregation methods as axes of comparison for the models, each representing a different approach to model aggregation: the seminal FedAvg (McMahan et al., 2017) algorithm, FedProx (Li et al., 2020a), which adds a proximal loss term to mitigate drift between local and global weights, SCAFFOLD (Karimireddy et al., 2020), which corrects client drift by using control variates to steer local updates towards the global model, FedAdam (Reddi et al., 2021), which decouples server-side and client-side optimization and employs the Adam optimizer (Kingma & Ba, 2017) at the server for model aggregation, FedBN (Li et al., 2021) which accommodates data heterogeneity by allowing clients to maintain their personal batch statistics, and by construction is only applicable to models with BN layers, FixBN (Zhong et al., 2024), which fixes the batch statistics halfway through training in order to combat some of BN’s issues, and FedPer (Arivazhagan et al., 2019) which personalizes the FL process by keeping the weights of the classifier head private to each client. We note our proposal is an architectural one which is aggregation method-agnostic, thus we selected these widely known aggregation methods to represent a spectrum of strategies, from standard averaging to methods addressing client drift and personalization. This provides a robust comparison concentrated on the model architectures.
@@
Evaluation metrics. For Fed-ISIC2019, we report the average balanced accuracy due to heavy class-imbalance as in (Ogier du Terrail et al., 2022). For FedChest, a multi-label classification task with imbalanced classes, we report the mean AUROC on the held-out test in this section and more metrics in Appendix C.2. We report the average accuracy for the other 3 datasets. In pFL settings, the objective is providing good in-federation models so we report the average metrics of the best local models, as suggested in (Zhang et al., 2023).
@@
Implementation details. We select hyper-parameters for each dataset by tuning the BN-ResNet (using the ranges detailed in Appendix A.2) and then use the same parameters for all models. This means the results in Section 4.2 are a conservative floor of the improvements that can be achieved, and tuning for ANFR can further increase improvements, as we will show in the next section. In Fed-ISIC2019 clients use Adam with a learning rate of 5e-4 and a batch size of to train for rounds of steps. This setup is distinct from the one used in Ogier du Terrail et al. (2022) resulting in performance improvements for all models. In Appendix B.1 we provide additional results using the original settings. In FedChest clients use Adam with a learning rate of 5e-4 and a batch size of to train for rounds of steps. For DP-training in FedPathology, we set the probability of information leakage to , as is common, the noise multiplier to , the gradient max norm to , and train for rounds, which is the point where the models have expended a privacy budget of . For CelebA and CIFAR-10 we follow the settings of Qu et al. (2022); Pieri et al. (2023) which were tuned by the authors. All experiments are run in a simulated FL environment with NVFLARE (Roth et al., 2022) and PyTorch (Paszke et al., 2019), using 2 NVIDIA A100 GPUs for training. We report the mean and standard deviation across 3 seeds.
@@ 4.2 Performance Analysis and Comparison
GFL scenario. Average results for all datasets, models, and GFL aggregation methods are presented in Table 2. First, we observe that GN does not consistently outperform the vanilla ResNet, supporting our pursuit of a more reliable alternative. For instance, GN is outperformed by BN in half of the tested aggregation methods on Fed-ISIC2019 and FedChest. Second, the sub-optimality of CA operating on BN-normalized features is evident, as the SE model frequently performs worse than BN-ResNet, notably across all aggregation methods on FedChest. NF-ResNet shows strong performance across all tasks and methods, confirming the potential of replacing activation normalization with weight standardization in FL. However, our proposed ANFR model consistently outperforms NF-ResNet, often by a considerable margin. For example, on Fed-ISIC2019 with SCAFFOLD, ANFR surpasses NF-ResNet’s mean balanced accuracy by more than 3%. For the FedChest dataset, we employ a large batch size of 128 to maximize the probability that all classes are represented in each batch, following best practices for multi-label, class-imbalanced datasets. This is further analyzed in a batch size ablation in Appendix C.3. ANFR emerges as the top-performing model across aggregation methods and our results indicate that integrating CA with SWS networks provides significant performance gains, suggesting that channel attention is a crucial component in designing effective FL models.
@@
pFL scenario. Table 3 presents the results for pFL scenarios on Fed-ISIC2019 and FedChest. In FedChest, where images are grayscale and we use a large batch size, FedBN and FedPer are virtually equal: BN-ResNet achieves an AUROC of 83.38% with FedBN and 83.39% with FedPer, indicating that the estimated BN statistics closely match the true ones. GN-ResNet attains 83.73% with FedPer, slightly outperforming BN-ResNet, but ANFR with FedPer is the most performant option across both aggregation methods, yielding a mean AUROC of 83.8%. Conversely, under the severe label and quantity skew on Fed-ISIC2019, employing FedBN improves performance over FedPer for models employing BN. ANFR achieves the highest balanced accuracy of 84.94% nonetheless. Notably, GN performs worse than BN on Fed-ISIC2019, and the ineffectiveness of combining BN and CA is further evidenced, as SE-ResNet is outperformed by BN-ResNet in all scenarios. These findings demonstrate that adopting ANFR enhances performance across both datasets, leading to the best overall models. Unlike the trade-offs observed with BN-FedBN and GN-FedPer combinations, ANFR consistently outperforms other architectures across varying levels of data heterogeneity.
Tuning in favor of ANFR in Fed-ISIC2019. As noted in Section 4.1 and further detailed in Appendix A.2, our hyper-parameters are chosen after tuning the baseline BN-ResNet and not ANFR, meaning the reported improvement in the results tables is a conservative floor of the improvement that can be achieved. To illustrate the real impact of our approach, we double the number of local steps in Fed-ISIC2019, keeping all other settings constant. As seen in Table 4, the performance of ANFR increases by compared to Table 2, while its improvement over the best baseline becomes twice as big. While this experimental setting favors ANFR, the performance of BN-ResNet is now lower, so this is not the setting we report in Table 2. The same methodology has been applied for all experimental settings. Despite optimizing for the baselines, ANFR still remains the best option, which greatly bolsters how exciting our results are.
BN-ResNet | GN-ResNet | SE-ResNet | NF-ResNet | ANFR | |
FedAvg | 64.52 | 66.16 | 67.55 | 71.76 | 76.34 |
4.3 Cross-device experiments on CelebA
@@
Table 5 presents the results of our models on the cross-device setting of CelebA, which contains 200,288 samples across 9,343 clients. While the binary classification task is relatively straightforward for individual clients, it poses challenges at the server level due to the vast number of clients and significant quantity and class skews—some clients have only a few samples or labels from a single class. We observe that ANFR outperforms the baseline models, demonstrating its adaptability across diverse FL scenarios.
@@ 4.4 Sample-level Differentially Private Training
In privacy-preserving scenarios involving differential privacy (DP), BN cannot be used as calculating mini-batch statistics violates privacy-preservation so it is customarily replaced by GN. We demonstrate the utility of ANFR in such settings using the FedPathology setup described in Section 4.1. We train using DP-SGD with strict sample-level privacy guarantees: following good practices, the probability of information leakage is set to , the noise multiplier is set to 1.1 and the gradient max norm to 1. We employ a privacy budget of , followed by training without privacy constraints (), to illustrate the privacy/utility trade-off of each model. From the results presented in Table 6, we observe that with an unrestricted privacy budget, GN and ANFR perform comparably. However, when a strict budget is enforced GN suffers a sharp performance decrease of 17%, as expected following previous research (Klause et al., 2022), whereas ANFR’s average accuracy is reduced by only 3%. ANFR’s robustness under DP may be attributed to its reliance on weight standardization, which has been shown to benefit from additional regularization (Brock et al., 2021b; Zhuang & Lyu, 2024) such as that provided by DP-SGD’s gradient clipping and gradient noising. Our experiments show DP training induces a regularization effect that disproportionately benefits NF models like ANFR, an observation also reported by De et al. (2022). These findings make ANFR a promising candidate for furthering development and adoption of DP training in FL, thereby enhancing the privacy of source data contributors, such as patients.
@@ 4.5 Attention Mechanism Comparison
Next, we investigate the impact of different attention mechanisms on performance. We compare the SE module used in previous sections with ECA (Wang et al., 2020), and CBAM (Woo et al., 2018). ECA replaces SE’s fully-connected layers with a more efficient 1-D convolution to capture local cross-channel interactions. CBAM combines channel and spatial attention and utilizes both max and average pooling to extract channel representations. From Table 7 we observe that even the lowest-performing module on each dataset outperforms all baseline models from Tables 2 and 5, proving the robustness of our approach. No single mechanism consistently performs best, making further exploration of attention modules an interesting avenue for future work.
@@
4.6 Computational Overhead
Additionally, to gauge the computational overhead of ANFR, and by extension its applicability in low-resource environments, we compare training times for BN-ResNet-26 with those for ANFR-26 using ECA as the attention mechanism. The batch size is set to 32, and we measure the average time per iteration of forward + backward pass across 100 iterations using PyTorch’s profiler. We do this for two distinct scenarios: devices without a CUDA-enabled GPU (e.g., smartphones), and devices with CUDA-enabled GPUs (e.g., edge devices such as Nvidia Jetson). The results in Table 8 show ANFR introduces marginal overhead ( without CUDA, with CUDA) while providing a significant performance advantage, showcasing its practicality in resource-constrained settings.
Scenario | Without CUDA | With CUDA | |||
Metric | Forward | Backward | Total | CPU time | GPU time |
BN-ResNet-26 | 297ms | 672ms | 969ms | 12ms | 22ms |
ANFR-26 (ECA) | 353ms | 717ms | 1s 70ms | 9ms | 26ms |
4.7 Qualitative Localization Performance
We assess the localization capability of each architecture after FL training with the best aggregation method on FedChest, SCAFFOLD. We compare the bounding box annotations provided by Wang et al. (2017) with Grad-CAM++ (Chattopadhay et al., 2018) heatmaps generated for samples labeled Atelectasis or Pneumonia from the FedChest test set. Figure 4 shows that ANFR’s heatmaps more closely align with the annotated bounding boxes. This improved localization aids model interpretability, which is crucial in areas like medical imaging.
4.8 Comparing Different Architectural Families
Finally, we relax the constraint of comparing between models of similar computational overhead and structure by broadening our evaluation to include models from different families. We compare ResNet-50 and ANFR-50 with EfficientNetv2-Small (Tan & Le, 2021), ViT-Small (Dosovitskiy et al., 2020; Wightman, 2019), ConvNeXt-Small (Liu et al., 2022) and FedConv (Xu et al., 2024). We compare, using FedAvg as the aggregation method, on the bases of accuracy, theoretical overhead, and practical overhead to provide more informative insights. The results, presented in Table 9, show that care is needed when choosing a network in order to balance performance and computational demands. This is especially true in cross-device FL scenarios, where straggler handling and limited compute concerns might render prohibitive the cost of otherwise performant networks like ConvNeXt (2 memory) or FedConv (2.64 latency). ANFR is shown to be more memory-efficient than a vanilla ResNet (due to removing BN), and occurring minimal overhead in terms of latency while being a top competitor in terms of accuracy.
Accuracy/AUROC(∗) | Theoretical Overhead | Measured Overhead | |||||
Architecture | FedChest∗ | Fed-ISIC2019 | CIFAR-10 | Params | GFLOPs | Memory in GB | Latency in ms |
ResNet-50 | 82.8 | 66.01 | 91.71 | 23.52 | 4.09 | 4.01 (1) | 32.38 (1) |
EfficientNetV2S | 79.53 | 69.24 | 93.93 | 20.19 | 2.85 | 5.91 (1.47) | 79.98 (2.47) |
ConvNeXt-Small | 82.28 | 77.71 | 97.73 | 49.46 | 8.68 | 8.16 (2.03) | 53.32 (1.65) |
ViT-Small | 83.53 | 73.21 | 97.48 | 21.68 | 4.96 | 3.36 (0.84) | 23.26 (0.71) |
FedConv-InvertUp | 82.68 | 78.32 | 97.26 | 25.62 | 4.61 | 5.35 (1.33) | 85.46 (2.64) |
ANFR-50 w/ ECA | 83.62 | 75.07 | 97.13 | 23.52 | 4.09 | 3.72 (0.93) | 40.16 (1.24) |
ANFR-50 w/ SE | 83.49 | 74.78 | 97.42 | 26.06 | 4.09 | 3.86 (0.96) | 44.67 (1.38) |